# On Minimal Transversals of Maximal Cliques in Graphs

### v1

@@ Abstract
A hypergraph is conformal if it is the family of maximal cliques of a graph. In this paper we are interested in the problem of determining when is the family of minimal transversal of maximal cliques of a graph conformal. Such graphs are called clique dually conformal (CDC for short). As our main results, we completely characterize CDC graphs within the families of triangle-free graphs and split graphs. Both characterizations lead to polynomial-time recognition algorithms. We also show that the class of CDC graphs is closed under substitution, in the strong sense that substituting a graph $H$ for a vertex of a graph $G$ results in a CDC graph if and only if both $G$ and $H$ are CDC.
@@ 1 Introduction
A hypergraph is a finite set of finite sets called hyperedges. A hypergraph is said to be Sperner [44] (also called simple [6, 7] or a clutter [41]) if no hyperedge contains another, and conformal if any set of vertices such that any two belong to a hyperedge is itself contained in a hyperedge (see, e.g., [41]). Sperner hypergraphs and conformal hypergraphs have been extensively studied in the literature, due to their numerous applications in combinatorics and in many other fields of mathematics and computer science (see, e.g., [1, 20, 5, 6]). Sperner hypergraphs enjoy a useful duality relation via the operation mapping a Sperner hypergraph $\mathcal{H}$ to its dual hypergraph $\mathcal{H}^{d}$ (also called the blocker of $\mathcal{H}$; see, e.g., Schrijver [41]), defined as the collection of all minimal transversals (also called minimal hitting sets), that is, inclusion-wise minimal sets of vertices intersecting each hyperedge in at least one vertex. The useful duality relation states that $\mathcal{H}^{dd}=\mathcal{H}$, that is, when restricted to the family of Sperner hypergraphs, the duality operator is an involution (see, e.g., Berge [6], Schrijver [41], and Crama and Hammer [17]). A similar duality holds for conformal hypergraphs, for the operator of mapping a conformal hypergraph $\mathcal{H}$ to its antiblocker $\mathcal{H}^{a}$, defined as the set of all inclusion-wise maximal sets of vertices intersecting each hyperedge in at most one vertex. If $\mathcal{H}$ is conformal, then $\mathcal{H}^{aa}=\mathcal{H}$ (as shown by Woodall [47, 48]; see also Schrijver [41]).
@@
The close connections with graphs stem from the fact that hypergraphs that are both conformal and Sperner are precisely the collections of maximal cliques of graphs (see [5]). More precisely, for every conformal Sperner hypergraph $\mathcal{H}$, there exists a graph $G$ such that $\mathcal{H}$ is the clique hypergraph $\mathcal{C}(G)$ of $G$, the hyperedges being exactly the maximal cliques of $G$. For example, using this connection, the fact that $\mathcal{H}^{aa}=\mathcal{H}$ when restricted to conformal hypergraphs is a simple consequence of the fact that graph complementation operation is an involution. Furthermore, exploiting the connection with graphs, the approach from [12] was shown to have applications in algorithmic graph theory, leading to a polynomial-time algorithm for checking, for any fixed positive integer $k$, if the upper clique transversal number of a given graph $G$ is at most $k$. The upper clique transversal number of a graph is defined as the maximum cardinality of a minimal transversal of maximal cliques; we refer to the recent work of Milanič and Uno [39] for more details.
@@ 2 Preliminaries
Graphs. All graphs considered in this paper are finite, simple, and undirected, except for Section 8, where we also consider directed graphs. The neighborhood of a vertex $v\in V(G)$, that is, the set of vertices adjacent to $v$ in $G$, is denoted by $N_{G}(v)$. The closed neighborhood of $v$ is denoted by $N_{G}[v]$ and defined as $N_{G}(v)\cup\{v\}$. Given a set $S\subseteq V(G)$, the neighborhood of $S$ is denoted b $N_{G}(S)$ and defined as the set of vertices in $V(G)\setminus S$ that have a neighbor in $S$. In all these notations, the subscript $G$ is omitted when the graph is clear from context. A clique in a graph is a set of pairwise adjacent vertices, an independent set (also called a stable set) is a set of pairwise non-adjacent vertices, a vertex cover is a set of vertices intersecting all edges, a matching is a set of pairwise disjoint edges, and a matching is perfect if every vertex belongs to a matching edge. A clique (resp., independent set) is maximal if it is not contained in any larger clique (resp., independent set). A clique transversal in a graph is a set of vertices containing at least one vertex from each maximal clique; a clique transversal is minimal if it does not contain any smaller clique transversal. Given a graph $G$, its complement $\overline{G}$ is defined by the same vertex set, $V(\overline{G})=V(G)$, and the complementary edge set: two distinct vertices $u,v\in V(G)$ are adjacent in $\overline{G}$ if and only if they are nonadjacent in $G$. (Recall that we restrict ourselves to simple graphs.) A graph is triangle-free if it does not have a clique of size three, split if its vertex set can be partitioned into a clique and an independent set, and cobipartite if its complement is bipartite, or, equivalently, its vertex set is a union of two cliques. We denote by $\cong$ the graph isomorphism relation.
@@ 3.1 Definition, basic properties, and examples
Since $G$ is a CDC graph, the clique hypergraph $\mathcal{C}(G)$ is dually conformal, implying, by Theorem 2.2, that $\mathcal{C}^{d}(G)=\mathcal{C}(G^{c})$. Since the clique hypergraphs are Sperner, the above equation also implies that $\mathcal{C}(G)=\mathcal{C}^{d}(G^{c})$. Hence, the dual of the hypergraph $\mathcal{C}(G^{c})$ is conformal, showing that $G^{c}$ is a CDC graph. The above equation implies that the minimal clique transversals of $G^{c}$ are exactly the maximal cliques of $G$. Consequently, we have $G^{cc}=G$. ∎
@@ 3.2 Graphs for which $G^{c}$ coincides with either $G$ or $\overline{G}$
Let $G$ be a graph that is not complete. Then, $G$ contains a vertex $u$ that is not universal. Let $S=V(G)\setminus N(u)$, that is, $S$ is the set consisting of $u$ and all non-neighbors of $u$ in $G$. Then $S$ intersects all maximal cliques in $G$ because any maximal clique in $G$ that does not contain any non-neighbor of $u$ must contain $u$, otherwise it would not be maximal. Since $S$ is a clique transversal in $G$, there exists a minimal clique transversal $T\subseteq S$. Any maximal clique $C$ containing $u$ does not contain any non-neighbors of $u$, hence $S\cap C=\{u\}$ and consequently $T\cap C=\{u\}$; in particular, this shows that $u\in T$. Fix a non-neighbor $w$ of $u$ and a maximal clique $D$ containing $w$. Then, the set $T\cap D$ is non-empty. Let $z$ be a vertex in $T\cap D$. Note that $u\not\in D$ and thus $z\neq u$. Furthermore, since $T\cap N(u)=\emptyset$, we infer that $z$ is a non-neighbor of $u$. Therefore, $T$ contains a pair of non-adjacent vertices $u$ and $z$, and hence is not a clique. ∎
@@
We translate this result into hypergraph terms. To this end, we first prove the following claim.
@@ 6.2 A characterization of twin-free triangle-free CDC graphs
In Corollary 6.16, the assumption that $G$ is twin-free is necessary. A triangle-free CDC graph that is not Kőnig-Egerváry can be obtained as follows. First, let $H=C_{5}\odot K_{1}$ be the corona of the $5$-cycle. Then $H$ is a triangle-free CDC graph, by Corollary 6.7. Let $G$ be the graph obtained from $H$ by substituting $2K_{1}$ into each vertex of degree $3$ in $H$. Clearly, $G$ is triangle-free, and by Theorem 5.5, $G$ is a CDC graph. However, $G$ is not Kőnig-Egerváry. The graph has $15$ vertices, hence its matching number is at most $7$ (in fact, it is exactly $7$). Its independence number is equal to the weighted independence number of the graph $H$ in which each vertex of the $5$-cycle has weight $2$ and each pendant vertex has weight $1$. Let $I$ be an independent set in $H$ and let $k$ be the number of vertices that the set contains from the $5$-cycle. Then $k\in\{0,1,2\}$ and the weight of $I$ is at most $5+k$, since the vertices from the $5$-cycle contribute a weight of $2k$ and the vertices outside the cycle contribute $5-k$. Thus, the independence number of $G$ is at most $7$ (in fact, it is exactly $7$). Consequently, the vertex cover number of $G$ is at least $15-7=8$, and since the matching number of $G$ is at most $7$, we conclude that $G$ is not Kőnig-Egerváry.
@@ 8 A relaxation of the CDC property: cycles of hypergraphs
We conclude the paper with a generalization of the concept of CDC graphs, or, more precisely, of pairs of CDC graphs and their clique-duals.
@@
Given a hypergraph $\mathcal{H}$, consider its co-occurrence graph $G(\mathcal{H})$ and denote by $\mathcal{H}^{c}$ the clique hypergraph of $G(\mathcal{H})$. By definition, for any $\mathcal{H}$ its conformalization $\mathcal{H}^{c}$ is Sperner, conformal and has the same vertex set as $\mathcal{H}$, $V(\mathcal{H}^{c})=V(\mathcal{H})$. Furthermore, $\mathcal{H}^{c}=\mathcal{H}$ if and only if $\mathcal{H}$ is Sperner and conformal. Note that in this case both operations $c$ and $d$ are involutions, that is, $\mathcal{H}^{cc}=\mathcal{H}^{dd}=\mathcal{H}$.
@@ Acknowledgements
The authors are grateful to Clément Dallard for helpful discussions. Part of the work for this paper was done in the framework of bilateral projects between Slovenia and the USA, partially financed by the Slovenian Research and Innovation Agency (BI-US-22/24/003, BI-US-22/24/076, BI-US/22–24–093, and BI-US/22–24–149). The work of the third author is supported in part by the Slovenian Research and Innovation Agency (I0-0035, research program P1-0285 and research projects J1-3001, J1-3002, J1-3003, J1-4008, J1-4084, and N1-0160) and by the research program CogniCom (0013103) at the University of Primorska. The second and fourth authors were working within the framework of the HSE University Basic Research Program. The work of the fifth author is partially supported by JSPS KAKENHI Grant Number JP17K00017, 20H05964 and 21K11757, Japan.

### v2

# Conformality of Minimal Transversals of Maximal Cliques

@@ Abstract
Given a hypergraph $\mathcal{H}$, the dual hypergraph of $\mathcal{H}$ is the hypergraph of all minimal transversals of $\mathcal{H}$. A hypergraph is conformal if it is the family of maximal cliques of a graph. In a recent work, Boros, Gurvich, Milanič, and Uno (Journal of Graph Theory, 2025) studied conformality of dual hypergraphs and proved several results related to this property, leading in particular to a polynomial-time algorithm for recognizing graphs in which all minimal transversals of maximal cliques have size at most $k$, for any fixed $k$. In this follow-up work, we provide a novel aspect to the study of graph clique transversals, by considering the dual conformality property from the perspective of graphs. More precisely, we study graphs for which the family of minimal transversals of maximal cliques is conformal. Such graphs are called clique dually conformal (CDC for short). It turns out that the class of CDC graphs is a rich generalization of the class of $P_{4}$-free graphs. As our main results, we completely characterize CDC graphs within the families of triangle-free graphs and split graphs. Both characterizations lead to polynomial-time recognition algorithms. Generalizing the fact that every $P_{4}$-free graph is CDC, we also show that the class of CDC graphs is closed under substitution, in the strong sense that substituting a graph $H$ for a vertex of a graph $G$ results in a CDC graph if and only if both $G$ and $H$ are CDC.
@@ 1 Introduction
A hypergraph is a finite set of finite sets called hyperedges (see Section 2 for more details). A hypergraph is said to be Sperner [44] (also called simple [6, 7] or a clutter [41]) if no hyperedge contains another, and conformal if any set of vertices such that any two belong to a hyperedge is itself contained in a hyperedge (see, e.g., [41]). Sperner hypergraphs and conformal hypergraphs have been extensively studied in the literature, due to their numerous applications in combinatorics and in many other fields of mathematics and computer science (see, e.g., [1, 20, 5, 6]). Sperner hypergraphs enjoy a useful duality relation via the operation mapping a Sperner hypergraph $\mathcal{H}$ to its dual hypergraph $\mathcal{H}^{d}$ (also called the blocker of $\mathcal{H}$; see, e.g., Schrijver [41]), defined as the collection of all minimal transversals (also called minimal hitting sets), that is, inclusion-wise minimal sets of vertices intersecting each hyperedge in at least one vertex. The useful duality relation states that $\mathcal{H}^{dd}=\mathcal{H}$, that is, when restricted to the family of Sperner hypergraphs, the duality operator is an involution (see, e.g., Berge [6], Schrijver [41], and Crama and Hammer [17]). A similar duality holds for Sperner hypergraphs that are also conformal, for the operator of mapping a conformal Sperner hypergraph $\mathcal{H}$ to its antiblocker $\mathcal{H}^{a}$, defined as the set of all inclusion-wise maximal sets of vertices intersecting each hyperedge in at most one vertex. If $\mathcal{H}$ is conformal and Sperner, then $\mathcal{H}^{aa}=\mathcal{H}$ (as shown by Woodall [47, 48]; see also Schrijver [41]).
@@
The close connections with graphs stem from the fact that hypergraphs that are both conformal and Sperner are precisely the collections of maximal cliques of graphs (see [5]). More precisely, for every conformal Sperner hypergraph $\mathcal{H}$, there exists a graph $G$ such that $\mathcal{H}$ is the clique hypergraph $\mathcal{C}(G)$ of $G$, the hyperedges being exactly the maximal cliques of $G$. For example, using this connection, the fact that $\mathcal{H}^{aa}=\mathcal{H}$ when restricted to conformal Sperner hypergraphs is a simple consequence of the fact that graph complementation operation is an involution. Furthermore, exploiting the connection with graphs, the approach from [12] was shown to have applications in algorithmic graph theory, leading to a polynomial-time algorithm for checking, for any fixed positive integer $k$, if the upper clique transversal number of a given graph $G$ is at most $k$. The upper clique transversal number of a graph is defined as the maximum cardinality of a minimal transversal of maximal cliques; we refer to the recent work of Milanič and Uno [39] for more details.
@@ 2 Preliminaries
Graphs. All graphs considered in this paper are finite, simple, and undirected, except for Section 8, where we also consider directed graphs. The neighborhood of a vertex $v\in V(G)$, that is, the set of vertices adjacent to $v$ in $G$, is denoted by $N_{G}(v)$. The closed neighborhood of $v$ is denoted by $N_{G}[v]$ and defined as $N_{G}(v)\cup\{v\}$. Given a set $S\subseteq V(G)$, the neighborhood of $S$ is denoted by $N_{G}(S)$ and defined as the set of vertices in $V(G)\setminus S$ that have a neighbor in $S$. In all these notations, the subscript $G$ is omitted when the graph is clear from context. A clique in a graph is a set of pairwise adjacent vertices, an independent set (also called a stable set) is a set of pairwise non-adjacent vertices, a vertex cover is a set of vertices intersecting all edges, a matching is a set of pairwise disjoint edges, and a matching is perfect if every vertex belongs to a matching edge. A clique (resp., independent set) is maximal if it is not contained in any larger clique (resp., independent set). A clique transversal in a graph is a set of vertices containing at least one vertex from each maximal clique; a clique transversal is minimal if it does not contain any smaller clique transversal. Given a graph $G$, its complement $\overline{G}$ is defined by the same vertex set, $V(\overline{G})=V(G)$, and the complementary edge set: two distinct vertices $u,v\in V(G)$ are adjacent in $\overline{G}$ if and only if they are nonadjacent in $G$. (Recall that we restrict ourselves to simple graphs.) A graph is triangle-free if it does not have a clique of size three, split if its vertex set can be partitioned into a clique and an independent set, and cobipartite if its complement is bipartite, or, equivalently, its vertex set is a union of two cliques. We denote by $\cong$ the graph isomorphism relation.
@@ 3.1 Definition, basic properties, and examples
Since $G$ is a CDC graph, the clique hypergraph $\mathcal{C}(G)$ is dually conformal, implying, by Theorem 2.2, that $\mathcal{C}^{d}(G)=\mathcal{C}(G^{c})$. Since the clique hypergraphs are Sperner, the above equation and the fact that $\mathcal{H}^{dd}=\mathcal{H}$ for every Sperner hypergraph $\mathcal{H}$ (see, e.g., [6, 41, 17]), imply that $\mathcal{C}(G)=(\mathcal{C}^{d}(G))^{d}=(\mathcal{C}(G^{c}))^{d}=\mathcal{C}^{d}(G^{c})$. Hence, the dual of the hypergraph $\mathcal{C}(G^{c})$ is conformal, showing that $G^{c}$ is a CDC graph. The above equation implies that the minimal clique transversals of $G^{c}$ are exactly the maximal cliques of $G$. Consequently, we have $G^{cc}=G$. ∎
@@ 3.2 Graphs for which $G^{c}$ coincides with either $G$ or $\overline{G}$
Let $G$ be a graph that is not complete. Then, $G$ contains a vertex $u$ that is not universal. Let $S=V(G)\setminus N(u)$, that is, $S$ is the set consisting of $u$ and all non-neighbors of $u$ in $G$. Then $S$ intersects all maximal cliques in $G$ because any maximal clique in $G$ that does not contain any non-neighbor of $u$ must contain $u$; otherwise it would not be maximal. Since $S$ is a clique transversal in $G$, there exists a minimal clique transversal $T\subseteq S$. Any maximal clique $C$ containing $u$ does not contain any non-neighbors of $u$. Hence, $S\cap C=\{u\}$ and consequently $T\cap C=\{u\}$; in particular, this shows that $u\in T$. Fix a non-neighbor $w$ of $u$ and a maximal clique $D$ containing $w$. Then, the set $T\cap D$ is non-empty. Let $z$ be a vertex in $T\cap D$. Note that $u\not\in D$ and thus $z\neq u$. Furthermore, since $T\cap N(u)=\emptyset$, we infer that $z$ is a non-neighbor of $u$. Therefore, $T$ contains a pair of non-adjacent vertices $u$ and $z$, and hence is not a clique. ∎
@@
We now prove an analogous result for hypergraphs. Given a hypergraph $\mathcal{H}$, consider its co-occurrence graph $G(\mathcal{H})$ and denote by $\mathcal{H}^{c}$ the clique hypergraph of $G(\mathcal{H})$. By definition, for any $\mathcal{H}$ its conformalization $\mathcal{H}^{c}$ is Sperner, conformal, and has the same vertex set as $\mathcal{H}$, $V(\mathcal{H}^{c})=V(\mathcal{H})$. Furthermore, $\mathcal{H}^{c}=\mathcal{H}$ if and only if $\mathcal{H}$ is Sperner and conformal. Note that in this case both operations $c$ and $d$ are involutions, that is, $\mathcal{H}^{cc}=\mathcal{H}^{dd}=\mathcal{H}$.
@@ 6.2 A characterization of twin-free triangle-free CDC graphs
In Corollary 6.16, the assumption that $G$ is twin-free is necessary. A triangle-free CDC graph that is not Kőnig-Egerváry can be obtained as follows. First, let $H=C_{5}\odot K_{1}$ be the corona of the $5$-cycle. Then $H$ is a triangle-free CDC graph, by Corollary 6.7. Let $G$ be the graph obtained from $H$ by substituting $2K_{1}$ into each vertex of degree $3$ in $H$. Clearly, $G$ is triangle-free, and by Theorem 5.5, $G$ is a CDC graph. However, $G$ is not Kőnig-Egerváry. The graph has $15$ vertices and, hence, its matching number is at most $7$ (in fact, it is exactly $7$). Its independence number is equal to the weighted independence number of the graph $H$ in which each vertex of the $5$-cycle has weight $2$ and each pendant vertex has weight $1$. Let $I$ be an independent set in $H$ and let $k$ be the number of vertices that the set contains from the $5$-cycle. Then $k\in\{0,1,2\}$ and the weight of $I$ is at most $5+k$, since the vertices from the $5$-cycle contribute a weight of $2k$ and the vertices outside the cycle contribute $5-k$. Thus, the independence number of $G$ is at most $7$ (in fact, it is exactly $7$). Consequently, the vertex cover number of $G$ is at least $15-7=8$, and since the matching number of $G$ is at most $7$, we conclude that $G$ is not Kőnig-Egerváry.
@@ 8 A relaxation of the CDC property: cycles of hypergraphs
We conclude the paper with a generalization of the concept of CDC graphs, or, more precisely, of pairs of CDC graphs and their clique-duals (see the paragraph after the proof of Proposition 8.3).
@@
Recall from Section 3.2 that the conformalization of a hypergraph $\mathcal{H}$ is the hypergraph denoted by $\mathcal{H}^{c}$ and defined as the clique hypergraph of the co-occurrence graph of $\mathcal{H}$.
@@ Acknowledgements
The authors are grateful to Clément Dallard for helpful discussions and to the two anonymous reviewers for their valuable suggestions. Part of the work for this paper was done in the framework of bilateral projects between Slovenia and the USA, partially financed by the Slovenian Research and Innovation Agency (BI-US-22/24/003, BI-US-22/24/076, BI-US/22–24–093, BI-US/22–24–149, and BI-US/24–26–088). The work of the third author is supported in part by the Slovenian Research and Innovation Agency (I0-0035, research program P1-0285 and research projects J1-3003, J1-4008, J1-4084, J1-60012, and N1-0370) and by the research program CogniCom (0013103) at the University of Primorska. The second and fourth authors were working within the framework of the HSE University Basic Research Program. The work of the fifth author is partially supported by JSPS KAKENHI Grant Number JP17K00017, 20H05964 and 21K11757, Japan.